

********************************************************************
UARS WEEKLY STATUS REPORT - 14 MAY 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began day only operations on Tuesday, May 11 (99131). Full
solar data sets were being taken during this time however no stellar
calibrations were made. Normal operations resumed on Friday, May 14
(99134).
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems.
***********
PEM
***********
PEM is active under PMON control. The FOT and the PEM team are
completing close-out documentation on this problem; however progress has
been hampered with the departure of Peter Thomas.
While under PMON control, the PEM instrument computes energy deposition
from the particle spectrometers.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on May 11 and is scheduled to operate through
May 28.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII has just been activated, and is operating normally.
Observations are planned to continue until May 30.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
10 days CLA (7.0) PROD (L1-L3) (22R-10C=12L)
40 days PEM (3.32) PROD (OA-L2) (753R-668C-1F=84L)
108 days WIN (5.11) PROD (LOG) (2014R-1538C=476L)
49 days WIN (5.11) PROD (L1-L3) (2014R-440-3F=1571L)
The CDHF will load March 1999 current day DAAC data in May 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 01 May 1999 (Orbit 41725, GMT Day 121) through 07 May
1999 (GMT Day 127).
Observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 124/1237 41777 On
HRDI 126/0137 41800 Full Time
The beta angle ranged between the angles of 2.1 degrees to 26.0 degrees for
this report period. The beta angle is now increasing toward a maximum
angle of 77 degrees on 21 May 1999 (DOY 141).
Extra ( eleven) SOLSTICE Solar Offset experiments were performed this week
(spanning days 121 through 126) at the request of SUSIM.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 30 April 1999 and 07 May 1999 are provided in Appendix
D.
The clock error ranged from -5.4 to 4.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 00
% HALOE 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/117 - 2287.496987
Transponder B: 99/118 - 2287.498739
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
aendix E.
The
data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 12 minutes, 37.108 seconds (an increase of 0.448 seconds since
the last report period). This is a 0.0362 percent data loss which equals a
99.9638 percent data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The start of end-to-end testing is scheduled for next week.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT attended tOC meeting
s on 03 and 07 May 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 11
June 1999 (DOY 162).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 May 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41777 124/1237 99-157 HALOE Power ON
41781 124/1854 99-158 HALOE Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 May 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 May 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMAN
= 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -72.8/-95.2 -3.68/-4.62 +0.0/ +0.0 16.6 0.000 0.0
2 +44.8/-5.6 +5.47/+4.51 +31.6/-14.4 24.2 1.014 21.4
3 +33.6/-67.2 +4.51/+3.23 +31.2/-15.2 24.0 1.016 21.0
07 May 1999 - GMT Day 127 Beta = 26.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -78.4/-100.8 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +39.2/-16.8 +6.11/+4.83 +30.4/-15.2 24.8 1.023 22.1
3 +22.4/-50.4 +5.15/+3.87 +30.4/-16.0 24.8 1.024 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 May 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-024 FOT 41741 122/0302 20809 Late Acquisition
99-025 FOT 41755 123/0132 14170 Generic Late Acquisition
99-023 FOT 41815 127/0148 14170 Generic Late Acquisition
99-023 FOT 41842 128/2048 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-024 FOT 41741 122/0302 20809 Late Acquisition
99-025 FOT see list May 99 14170 Generic Late Acq
SEVNTFW99/19